OpenStudy (anonymous):

what is the point-slope form of a line with slope 6 that contains the point (1,2)?

OpenStudy (anonymous):

(y-y2)=m(x-x2) Where m=slope

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

Welcome...what is your answer?

OpenStudy (anonymous):

y-2=6(x-1)

OpenStudy (astrophysics):

\[y-y _{1}=m(x-x _{1}) \] well here is the formula in LaTeX, for a better reference, and yeah it's right, but you may want to put it in y = mx+b form.

OpenStudy (anonymous):

good job